# A Multi-Level Superoptimizer for Tensor Programs

### v2

@@ Abstract.
We introduce Mirage, the first multi-level superoptimizer for tensor programs. A key idea in Mirage is $\mu$Graphs, a uniform representation of tensor programs at the kernel, thread block, and thread levels of the GPU compute hierarchy. $\mu$Graphs enable Mirage to discover novel optimizations that combine algebraic transformations, schedule transformations, and generation of new custom kernels. To navigate the large search space, Mirage introduces a pruning technique based on abstraction that significantly reduces the search space and provides a certain optimality guarantee. To ensure that the optimized $\mu$Graph is equivalent to the input program, Mirage introduces a probabilistic equivalence verification procedure with strong theoretical guarantees. Our evaluation shows that Mirage outperforms existing approaches by 1.1-2.9$\times$ even for DNNs that are widely used and heavily optimized. Mirage is publicly available at https://github.com/mirage-project/mirage.
1. Introduction
@@
Enabling high-performance execution of deep neural networks (DNNs) on GPUs is critical for modern ML applications. Today’s DNN frameworks generally specify DNN computation using tensor programs, which are directed acyclic graphs whose nodes and edges represent tensor algebra operators (e.g., matrix multiplications) and tensors (i.e., $n$-dimensional array) shared between operators.
@@
To optimize an input tensor program, existing frameworks (e.g., PyTorch (pytorch) and TensorFlow (Tensorflow)) use manually designed rules to map the tensor program to expert-written GPU kernels. These approaches generally require extensive engineering efforts to design and implement optimization rules, and may miss some optimization opportunities. To address these challenges, recent work introduced automated approaches to optimizing tensor programs by searching over a comprehensive space of program transformations and applying them based on their performance on target GPUs. These approaches generally fall into two categories.
@@
The first category of work, including Halide (halide), TVM (tvm), and Ansor (ansor), is motivated by the idea of algorithm and schedule separation [^1] In the schedule optimization literature, an algorithm describes what to compute in a kernel and a schedule specifies how to compute the kernel. introduced in Halide and optimizes the schedule of a tensor program while fixing the algorithm. For a given algorithm, these optimizers automatically generate performant kernels by searching for possible strategies to execute the kernel on the target hardware. However, due to the linear algebra nature of DNNs, a tensor program can be represented by a wide spectrum of mathematically equivalent algorithms, and existing schedule-based optimizers only consider kernels whose algorithms are manually specified by users, resulting in missed optimization opportunities.
@@
The second category of work, including TASO (TASO), Grappler (tensorflow_grappler), Tensat (yang2021equality), and PET (wang2021pet), considers algebraic transformations, which exploit mathematical equivalence between different algorithms for a tensor program. Example algebraic transformations include (1) converting a linear algebra operator into another such as a convolution to a matrix multiplication, (2) fusing multiple operators to reduce memory access and kernel overhead, and (3) reorganizing operators based on commutativity, associativity, and distributivity. These optimizers perform algebraic transformations at the algorithm level and require programmers to manually define the set of available kernels and their implementations. They are thus limited by the performance of the provided kernels.
@@
All existing automated optimization approaches, from both categories, still require programmers to manually specify a set of kernels (each defined by a tensor function), and then explore the search space of algebraic or schedule transformations. However, some advanced performance optimizations require coordinated transformations at the kernel, thread block, and thread levels of the GPU compute hierarchy, and involve introducing completely new kernel computations (e.g., a custom kernel that decomposes standard kernels and fuses only some of their computations). Such optimizations are not part of the search space of existing automated methods and must still be implemented manually.
@@
One such example is FlashAttention (dao2023flash) (see §8.2 for details), which optimizes attention (transformers) on GPUs by reordering operators at the algorithm level (algebraic transformations), reorganizing the computation across GPU kernels (yielding new custom kernels), and adapting the parallelization strategy of each kernel to the GPU architecture (schedule transformations). The transformations required for this example cannot be automatically discovered by existing frameworks and must therefore be implemented manually. An implementation of FlashAttention in Triton (tillet2019triton), a widely used tensor program optimizer, includes more than 700 lines of code (triton-flashattention).
@@
We present Mirage, the first multi-level superoptimizer for tensor programs. Mirage is able to automatically discover and verify sophisticated optimizations of tensor programs that require joint optimization of algebraic transformations, schedule transformations, and the discovery of new custom kernels.
@@
A key idea in Mirage is $\mu$Graphs, a hierarchical graph representation that specifies a tensor program at multiple levels of the GPU compute hierarchy. By uniformly treating the kernel, thread block, and thread levels, $\mu$Graphs can capture both algebraic and schedule transformations across these levels. Moreover, optimizing a $\mu$Graph can introduce new custom kernels, which go beyond both algebraic and schedule transformations. For example, Mirage automatically discovers the $\mu$Graphs that represent FlashAttention (dao2023flash) and its inference variant FlashDecoding (tri2023flashdecoding) as well as other $\mu$Graphs that outperform these manually designed kernels by up to 3.5$\times$ for certain use cases. Most of these Mirage-discovered optimizations are outside the search space of existing approaches.
Figure 1. An overview of Mirage.
@@
Figure 1 shows an overview of Mirage. Mirage first splits an input tensor program into subprograms that fall into the restricted Lax fragment. The Lax fragment, formally defined in §5, includes multi-linear operators such as matrix multiplication and convolution, division (which is useful for normalizations), and limited use of exponentiation (which is useful for activations). The partitioning into Lax subprograms reduces the optimization search space while preserving most optimization opportunities; it also enables Mirage’s probabilistic equivalence verifier.
@@
For each Lax subprogram, Mirage’s expression-guided generator uses an exhaustive search to find possible $\mu$Graphs equivalent to it. A key challenge Mirage has to address is its significantly larger search space compared to prior superoptimization techniques for ML. For example, TASO (TASO) and PET (wang2021pet) only search for tensor programs at the kernel level by using a fixed set of pre-defined kernels, while Mirage simultaneously considers superoptimization at the kernel, thread block, and thread levels. To efficiently navigate this significantly larger and more complex search space, Mirage introduces a novel pruning technique based on abstract expressions. This approach greatly reduces the number of $\mu$Graphs Mirage has to consider while providing a theoretical guarantee on the optimality of the discovered $\mu$Graphs (§4.3).
@@
For a $\mu$Graph discovered by Mirage, verifying its functional equivalence with the input program introduces another challenge, since the input and output tensors of a program include up to many millions of elements. A key idea behind Mirage is probabilistic equivalence verification, which performs random tests over finite fields to check equivalence between $\mu$Graphs. While random tests can hardly provide any correctness guarantees for general programs, Mirage relies on a novel theoretical result to show that the restrictions of the Lax fragment ensure that for Lax programs, random tests over finite fields offer strong correctness guarantees. We show that an algorithm for polynomial identity testing (PIT) (schwartz1980fast; zippel1979probabilistic) can be generalized to Lax programs, yielding a randomized algorithm for Lax program equivalence that can be made arbitrarily precise. Mirage uses this random algorithm to (probabilistically) ensure that each optimized program is equivalent to the input program.
@@
For each verified $\mu$Graph, Mirage’s $\mu$Graph optimizer maximizes its runtime performance by further considering potential tensor layouts, scheduling operators’ execution orders, and planning memory allocation at all of the kernel, thread block, and thread levels. Finally, Mirage returns an optimized tensor program based on the best discovered $\mu$Graph for each individual Lax subprogram.
@@
We evaluate Mirage on a variety of commonly used DNN benchmarks on NVIDIA A100 and H100 GPUs. Even for DNN benchmarks that are widely used and heavily optimized by existing systems such as the group-query attention used in LLMs (grattafiori2024llama3herdmodels), Mirage still outperforms current approaches by 1.1-2.9$\times$, by exploiting subtle custom kernels and optimizations missing in existing systems.
2. Multi-Level Graph Representation
@@
Figure 2. GPU compute and memory hierarchy.
@@
Each tensor program corresponds to one kernel graph, where each node represents a kernel running on an entire GPU and each edge is a tensor shared between kernels. All tensors in a kernel graph are stored in GPU device memory since different kernels cannot share data in register file or shared memory. Each node in a kernel graph can be a pre-defined kernel operator supported by existing kernel libraries such as convolution by cuDNN (cudnn) and matrix multiplication by cuBLAS (cublas). In addition, to enable fine-grained inter-kernel optimizations such as kernel fusion, a node in a kernel graph can also be a graph-defined kernel operator, whose semantic and behavior are defined by a lower-level (i.e., block) graph. As an example, the kernel operator in Figure 4b is a graph-defined operator specified by a block graph.
@@
A block graph specifies computation associated with a thread block [^3] In the CUDA programming model, a kernel’s computation is defined as computations for independent thread blocks., where each node denotes a block operator which specifies computation within a block and each edge (blue arrows in Figure 4b) is a tensor shared between block operators. Mirage saves all intermediate tensors within a block graph in GPU shared memory for two considerations. First, GPU shared memory offers much higher bandwidth than device memory, and this design allows Mirage to reduce device memory access by maximally saving intermediate results in shared memory. Second, for tensors whose sizes exceed shared memory capacity and must be stored in device memory, Mirage uses these tensors to split computation into multiple block graphs, each of which only contains tensors in shared memory. This separation does not introduce additional access to device memory.
Each block graph is also associated with a few properties to specify its execution, which we introduce as follows.
@@
Figure 3. Demonstrating how an input tensor is partitioned across blocks and for-loop iterations with imap and fmap.
@@
All blocks within a kernel are organized by a mesh with up to 3 dimensions, identified as $x$, $y$, and $z$. A block graph is associated with up to three grid dimensions that specify the number of blocks along the $x$, $y$, and $z$ dimensions. The block graph in Figure 4b launches $128$ blocks.
@@
First, for each input tensor to a graph-defined kernel operator (e.g., $X$, $G$, and $W$ in the kernel graph in Figure 4b), the associated block graph contains an imap, which specifies how the input tensor is partitioned into sub-tensors for individual blocks. For each grid dimension (i.e., $x$, $y$, or $z$), the imap maps it to (1) a data dimension of the input tensor or (2) a special replica dimension $\phi$. For (1), the mapped data dimension is equally partitioned across blocks along the grid dimension. For (2), the input tensor is replicated across these blocks. As an example, the block graph in Figure 4b takes three inputs — $\overline{X}$, $\overline{G}$, and $\overline{W}$, which represent the input tensors to each block. For $\overline{W}$, its $\mbox{\rm\em imap}=\{x\leftrightarrow d\}$ indicates that the $d$ dimension of tensor $W$ is partitioned into 128 equally sized chunks. As a result, $\overline{W}$ is of shape $[h=1024,d=32]$.
@@
Second, for each output tensor of a block graph (e.g., $\overline{Z}$ in Figure 4b), the block graph includes an omap, which specifies how the outputs of all blocks are concatenated to construct the final output of the kernel operator. In an omap, each grid dimension must map to a data dimension of the output tensor, since different blocks must save to disjoint tensors in device memory. For $\overline{Z}$ of shape $[b=16,d=32]$ in Figure 4b, its $\mbox{\rm\em omap}=\{x\leftrightarrow d\}$ indicates that blocks with the same $x$ index are concatenated along the $d$ dimension, resulting in a tensor $Z$ of shape $[b=16,d=4096]$.
For-loop dimensions.
@@
To fit large input tensors in shared memory and allow cache reuse, a second property associated with each block graph is for-loop dimensions, which collectively specify how many times the block graph is executed to complete a kernel. For example, the block graph in Figure 4b has a for-loop dimension $i=16$, indicating it is executed 16 times to finish the associated graph-defined kernel operator. Correspondingly, each input tensor to a block graph is first sent to an input iterator that loads a part of the tensor (e.g., $\overline{X}$, $\overline{G}$, and $\overline{W}$) from device memory to shared memory. Each input iterator is associated with an fmap to specify which part of the input tensor to load in each iteration. Formally, the fmap maps each for-loop dimension to (1) a data dimension of the input tensor or (2) the replica dimension $\phi$. Similar to the semantic of imap, the input tensor is equally partitioned along that dimension for (1) and replicated for (2). Figure 3 shows how an input matrix is partitioned across blocks and for-loop iterations with different imap and fmap.
In addition, a block graph contains for-loop accumulators (e.g., the two Accum operators in Figure 4b) to accumulate its output across iterations in shared memory. Similar to an input iterator, a for-loop accumulator is also associated with an fmap to specify how the output tensors of different iterations are combined to produce the accumulated results. Specifically, the fmap maps each for-loop dimension to either a data dimension, which results in concatenation of the output along that dimension, or the replica dimension $\phi$, which results in the output being accumulated in shared memory.
@@
A thread graph further reduces the scope of computation from a block to a single thread. Similar to a block graph, each thread graph is also associated with block dimensions, which specify the organization of threads within the block, and for-loop dimensions, which define the total number of iterations to finish the defined computation. Each thread graph includes input iterators, each of which loads an input tensor (e.g., $\overline{\overline{A}}$ and $\overline{\overline{B}}$ in Figure 4b) from GPU shared memory to register file, and output savers, each of which saves an output tensor from register file back to shared memory (e.g., $\overline{\overline{C}}$). A thread graph is the lowest level graph in a $\mu$Graph and only contains pre-defined thread operators.
@@
Each tensor in the kernel, block, or thread graph is associated with a tensor layout (omitted in Figure 4 for simplicity), which specifies how the tensor is linearized in memory. Note that tensor layout only affects the execution performance of a $\mu$Graph and has no impact on its output.
@@
Prior work separately considers algebraic (TASO; wang2021pet) or schedule transformations (halide; tvm; autohalide), while $\mu$Graphs can represent both in a uniform way. Specifically, the grid and for-loop dimensions and their corresponding mappings (i.e., imap, omap, and fmap) to tensor dimensions constitute a comprehensive search space of possible schedules for graph-defined operators. The hierarchical graphs at the kernel, block, and thread levels allow Mirage to explore algebraic transformations at these levels.
3. Case Study: RMSNorm
(a) Computation graph for RMSNorm and MatMul.
(b) The best $\mu$Graph discovered by Mirage.
Figure 4. Figure 4a is the computation graph for RMSNorm and MatMul. Figure 4b shows the best $\mu$Graph discovered by Mirage for computing RMSNorm and MatMul, which fuses the computation in a single kernel to reduce device memory access and kernel launch overhead, outperforms existing approaches by 1.9$\times$. Numbers in brackets indicate tensor shapes, and numbers in braces show the imap, omap, or fmap for the corresponding operators.
In this section, we use the root mean square layer normalization (RMSNorm) (zhang2019rootmeansquarelayer) as a case study to demonstrate the advantages of the $\mu$Graph representation and Mirage’s superoptimization approach. RMSNorm is a widely used normalization technique by recent large language models such as LLAMA-2-70B (llama2). Formally, RMSNorm takes two tensors $X$ and $G$ as inputs and regularizes the summed inputs of $X$ according to the root mean square static:
(1) $Y_{i}=\frac{X_{i}G_{i}}{\textrm{RMS}(X)},\textrm{RMS}(X)=\sqrt{\frac{1}{d}\sum_{j=1}^{d}X_{j}^{2}},$

### v3

# Mirage: A Multi-Level Superoptimizer for Tensor Programs

@@ Abstract
We introduce Mirage, the first multi-level superoptimizer for tensor programs. A key idea in Mirage is $\mu$Graphs, a uniform representation of tensor programs at the kernel, thread block, and thread levels of the GPU compute hierarchy. $\mu$Graphs enable Mirage to discover novel optimizations that combine algebraic transformations, schedule transformations, and generation of new custom kernels. To navigate the large search space, Mirage introduces a pruning technique based on abstraction that significantly reduces the search space and provides a certain optimality guarantee. To ensure that the optimized $\mu$Graph is equivalent to the input program, Mirage introduces a probabilistic equivalence verification procedure with strong theoretical guarantees. Our evaluation shows that Mirage significantly outperforms existing approaches even for DNNs that are widely used and heavily optimized. Mirage is publicly available at https://github.com/mirage-project/mirage.
1 Introduction
@@
Enabling high-performance execution of deep neural networks (DNNs) on GPUs is critical for modern ML applications. Today’s DNN frameworks generally specify DNN computation using tensor programs, which are directed acyclic graphs whose nodes and edges represent tensor algebra operators (e.g., matrix multiplication) and tensors (i.e., $n$-dimensional arrays) shared between operators.
@@
To optimize an input tensor program, existing frameworks (e.g., PyTorch [34] and TensorFlow [9]) use manually designed rules to map the tensor program to expert-written GPU kernels. These approaches generally require extensive engineering efforts to design and implement optimization rules, and they may miss certain optimization opportunities. To address these challenges, recent work has introduced automated approaches that optimize tensor programs by searching over a comprehensive space of program transformations and applying them based on their performance on target GPUs. These approaches generally fall into two categories.
@@
The first category of work, including Halide [35], TVM [13], and Ansor [51], is motivated by the idea of algorithm and schedule separation [^1] In the schedule optimization literature, an algorithm describes what to compute in a kernel and a schedule specifies how to compute the kernel. introduced in Halide and optimizes the schedule of a tensor program while fixing the algorithm. For a given algorithm, these optimizers automatically generate performant kernels by searching for possible strategies to execute the kernel on the target hardware. However, due to the linear algebra nature of DNNs, a tensor program can be represented by a wide spectrum of mathematically equivalent algorithms. Existing schedule-based optimizers only consider kernels whose algorithms are manually specified by users, resulting in missed optimization opportunities.
@@
The second category of work, including TASO, Grappler, Tensat, and PET, considers algebraic transformations, which exploit mathematical equivalence among different algorithms for a tensor program [25, 3, 48, 46]. Examples of algebraic transformations include (1) converting one linear algebra operator into another, such as transforming a convolution to a matrix multiplication; (2) fusing multiple operators to reduce memory access and kernel overhead; and (3) reorganizing operators based on commutativity, associativity, and distributivity. These optimizers perform algebraic transformations at the algorithm level and require programmers to manually specify the set of available operators and their implementations. They are thus limited by the performance of the provided kernels.
@@
All existing automated optimization approaches, from both categories, still require programmers to manually specify a set of kernels (each defined by a tensor function), and then explore the search space of algebraic or schedule transformations. However, some advanced performance optimizations require coordinated transformations across the kernel, thread block, and thread levels of the GPU compute hierarchy, and involve introducing completely new kernel computations (e.g., a custom kernel that decomposes standard kernels and fuses only certain computations). Such optimizations are not included in the search space of existing automated methods and must still be implemented manually.
@@
One such example is FlashAttention [17] (see §8.2 for details), which optimizes attention [47] on GPUs by reordering operators at the algorithm level (algebraic transformations), reorganizing the computation across GPU kernels (yielding new custom kernels), and adapting the parallelization strategy of each kernel to the GPU architecture (schedule transformations). The transformations required for this example cannot be automatically discovered by existing frameworks and must therefore be implemented manually. An implementation of FlashAttention in Triton [43], a widely used tensor program optimizer, contains more than 700 lines of code [8].
@@
We present Mirage, the first multi-level superoptimizer for tensor programs. Mirage automatically discovers and verifies sophisticated optimizations of tensor programs that require joint optimization of algebraic transformations, schedule transformations, and the discovery of new custom kernels.
@@
A key idea in Mirage is $\mu$Graphs, a hierarchical graph representation that specifies tensor programs across multiple levels of the GPU compute hierarchy. By uniformly treating the kernel, thread block, and thread levels, $\mu$Graphs can capture both algebraic and schedule transformations across these levels. Moreover, optimizing a $\mu$Graph can introduce new custom kernels, which go beyond both algebraic and schedule transformations. For example, Mirage automatically discovers the $\mu$Graphs representing FlashAttention [17] and its inference variant FlashDecoding [5] as well as other $\mu$Graphs that outperform these manually designed kernels by up to 2.2$\times$ for certain use cases. Most of these optimizations discovered by Mirage are outside the search space of existing methods.
Figure 1: An overview of Mirage.
@@
Figure 1 shows an overview of Mirage. Mirage first splits an input tensor program into subprograms that fall into the restricted Lax fragment. The Lax fragment, formally defined in §5, includes multi-linear operators such as matrix multiplication and convolution, division (useful for normalizations), and limited exponentiation (useful for activations). Partitioning a tensor program into Lax subprograms reduces the optimization search space while preserving most optimization opportunities; it also enables Mirage’s probabilistic equivalence verifier.
@@
For each Lax subprogram, Mirage’s expression-guided generator exhaustively searches for possible $\mu$Graphs equivalent to it. A key challenge Mirage must address is its significantly larger search space compared to prior superoptimization techniques. For example, TASO [25] and PET [46] search only for tensor programs at the kernel level, using a fixed set of pre-defined kernels, while Mirage considers superoptimization across the kernel, thread block, and thread levels. To efficiently navigate this significantly larger search space, Mirage introduces a novel pruning technique based on abstract expressions, which greatly reduces the number of $\mu$Graphs Mirage must consider while providing a certain theoretical guarantee on the optimality of the discovered $\mu$Graphs. Mirage further reduces the search space by focusing the search on the kernel and block levels and using a rule-based approach for the thread level.
@@
For a $\mu$Graph discovered by Mirage, verifying its functional equivalence with the input program introduces another challenge, since the input and output tensors of a program include up to many millions of elements. A key idea behind Mirage is probabilistic equivalence verification, which performs random tests over finite fields to check equivalence between $\mu$Graphs. While random tests typically provide limited correctness guarantees for general programs, Mirage leverages a novel theoretical result showing that the restrictions imposed by the Lax fragment ensure that, for Lax programs, random tests over finite fields offer strong correctness guarantees. Specifically, we show that a polynomial identity testing (PIT) algorithm [37, 54] can be generalized to Lax programs, yielding a randomized algorithm for Lax program equivalence that can be made arbitrarily precise. Mirage uses this randomized algorithm to (probabilistically) ensure that each optimized program is equivalent to the input program.
@@
For each verified $\mu$Graph, Mirage’s $\mu$Graph optimizer maximizes its runtime performance by further considering potential tensor layouts, scheduling operator execution orders, and planning memory allocation at all of the kernel, thread block, and thread levels. Finally, Mirage returns an optimized tensor program based on the best discovered $\mu$Graph for each individual Lax subprogram.
@@
We evaluate Mirage on a variety of commonly used DNN benchmarks on NVIDIA A100 and H100 GPUs. Even for DNN benchmarks that are widely used and heavily optimized by existing systems, such as the group-query attention used in LLMs [41], Mirage still outperforms current approaches by up to 3.3$\times$ by exploiting subtle custom kernels and optimizations missing in existing systems.
2 Multi-Level Graph Representation
@@
Figure 2: GPU compute and memory hierarchy.
@@
(a) Computation graph for RMSNorm and MatMul.
(b) The best $\mu$Graph discovered by Mirage.
Figure 3: Figure 3a is the computation graph for RMSNorm and MatMul. Figure 3b shows the best $\mu$Graph discovered by Mirage for computing RMSNorm and MatMul, which fuses the computation in a single kernel to reduce device memory access and kernel launch overhead, outperforms existing approaches by 1.9$\times$. Numbers in brackets indicate tensor shapes, and numbers in braces show the imap, omap, or fmap for the corresponding operators.
@@
Each tensor program corresponds to one kernel graph, where each node represents a kernel running on an entire GPU, and each edge is a tensor shared between kernels. All tensors in a kernel graph are stored in GPU device memory since different kernels cannot share data in register files or shared memory. Each node in a kernel graph can be a pre-defined kernel operator supported by existing kernel libraries such as convolution by cuDNN [15] and matrix multiplication by cuBLAS [16]. In addition, to enable fine-grained inter-kernel optimizations such as kernel fusion, a node in a kernel graph can also be a graph-defined kernel operator, whose semantic and behavior are defined by a lower-level (i.e., block) graph. As an example, the kernel operator in Figure 3b is a graph-defined operator specified by a block graph.
@@
A block graph specifies computation associated with a thread block [^3] In the CUDA programming model, a kernel’s computation is defined as computations for independent thread blocks., where each node denotes a block operator specifying computation within a block, and each edge (blue arrows in Figure 3b) is a tensor shared between block operators. Mirage stores all intermediate tensors within a block graph in GPU shared memory for two considerations. First, GPU shared memory offers much higher bandwidth than device memory, and this design allows Mirage to reduce device memory access by maximally saving intermediate results in shared memory. Second, for tensors whose sizes exceed shared memory capacity and must be stored in device memory, Mirage uses these tensors to split computation into multiple block graphs, each of which only contains tensors in shared memory. This separation does not introduce additional access to device memory.
Each block graph is also associated with properties specifying its execution, which we introduce below.
@@
Figure 4: Demonstrating how an input tensor is partitioned across blocks and for-loop iterations with imap and fmap.
@@
All blocks within a kernel are organized into a mesh with up to 3 dimensions, identified as $x$, $y$, and $z$. A block graph is associated with up to three grid dimensions that specify the number of blocks along the $x$, $y$, and $z$ dimensions. The block graph in Figure 3b launches $128$ blocks.
@@
First, for each input tensor to a graph-defined kernel operator (e.g., $X$, $G$, and $W$ in the kernel graph in Figure 3b), the associated block graph contains an imap, which specifies how the input tensor is partitioned into sub-tensors for individual blocks. For each grid dimension (i.e., $x$, $y$, or $z$), the imap maps it to either (1) a data dimension of the input tensor or (2) a special replica dimension $\phi$. For (1), the mapped data dimension is equally partitioned across blocks along the grid dimension. For (2), the input tensor is replicated across these blocks. For example, the block graph in Figure 3b takes three inputs—$\overline{X}$, $\overline{G}$, and $\overline{W}$—representing the input tensors to each block. For $\overline{W}$, its $\mbox{\rm\em imap}=\{x\leftrightarrow d\}$ indicates that the $d$ dimension of tensor $W$ is partitioned into 128 equally sized chunks. As a result, $\overline{W}$ has shape $[h=1024,d=32]$.
@@
Second, for each output tensor of a block graph (e.g., $\overline{Z}$ in Figure 3b), the block graph includes an omap, which specifies how the outputs of all blocks are concatenated to construct the final output of the kernel operator. In an omap, each grid dimension must map to a data dimension of the output tensor, since different blocks must store disjoint tensors in device memory. For $\overline{Z}$ with shape $[b=16,d=32]$ in Figure 3b, its $\mbox{\rm\em omap}=\{x\leftrightarrow d\}$ indicates that blocks with the same $x$ index are concatenated along the $d$ dimension, resulting in a tensor $Z$ with shape $[b=16,d=4096]$.
For-loop body.
@@
To fit large input tensors in shared memory and to overlap data loading from device memory with computation, a block graph can include a for-loop body, which is executed multiple times to complete a kernel. Often, the for loop in a kernel is followed by some post-processing. For example, when computing an average value, the for loop would perform the summation of $n$ values and the post-processing would divide by $n$. Mirage specifies the for-loop body of a block graph using input iterators, for-loop accumulators, and all operators in between, as shown in the orange box in Figure 3b). Each input tensor to a block graph first passes through an input iterator, which loads part of the tensor (e.g., $\overline{X}$, $\overline{G}$, and $\overline{W}$) from device memory into shared memory. Each input iterator is associated with an fmap to specify which part of the input tensor to load in each iteration. Formally, the fmap maps each for-loop dimension to either (1) a data dimension of the input tensor or (2) the replica dimension $\phi$. Similar to imap, the tensor is equally partitioned along that dimension for (1) and replicated for (2). Figure 4 shows how an input matrix is partitioned across blocks and for-loop iterations using different imap and fmap.
Each block graph is also associated with a for-loop dimension, which determines how many iterations the for-loop body is executed to complete the kernel. In addition, Mirage uses for-loop accumulators (e.g., the two Accum operators in Figure 3b) to accumulate intermediate results computed in each iteration (using standard accumulators, e.g., summation and max) and store the accumulated results in shared memory. Once the for-loop body is completed, Mirage proceeds to execute the remaining operators outside the for-loop body directly on the accumulated results. An output saver then saves the final result from shared memory back to device memory.
@@
A thread graph further reduces computation scope from a block to a single thread. Similar to a block graph, each thread graph is also associated with block dimensions, which specify the organization of threads within the block, and for-loop dimensions, which define the total number of iterations to finish the defined computation. Each thread graph includes input iterators, each of which loads an input tensor (e.g., $\overline{\overline{A}}$ and $\overline{\overline{B}}$ in Figure 3b) from shared memory into register files, and output savers, each of which stores an output tensor from register files back to shared memory (e.g., $\overline{\overline{C}}$). A thread graph is the lowest-level graph in a $\mu$Graph and contains only pre-defined thread operators.
@@
Each tensor in the kernel, block, or thread graph is associated with a tensor layout (omitted in Figure 3 for simplicity), specifying how the tensor is linearized in memory. Note that tensor layouts affect only the performance of a $\mu$Graph and have no impact on its output correctness.
Definition 2.1 ($\mu$Graph Validity).
A $\mu$Graph $G$ is valid if: (1) for each kernel, block, and thread operator $o\in G$, its input and output tensors match the specification of $o$; (2) all tensors in each kernel, block, and thread graph can reside in GPU device memory, shared memory, and register file, respectively; and (3) for each block and thread graph with a for-loop body, any path from an input to an output passes through exactly one input-iterator, one for-loop accumulator, and one output saver.
@@
Prior work separately considers algebraic [25, 46] or schedule transformations [35, 13, 31], while $\mu$Graphs can represent both in a uniform way. Specifically, the grid and for-loop dimensions and their corresponding mappings (i.e., imap, omap, and fmap) to tensor dimensions constitute a comprehensive search space of possible schedules for graph-defined operators. The hierarchical graphs across the kernel, block, and thread levels allow Mirage to explore algebraic transformations at these levels.
3 Case Study: RMSNorm
In this section, we use root mean square layer normalization (RMSNorm) [50] as a case study to demonstrate the advantages of the $\mu$Graph representation and Mirage’s superoptimization approach. RMSNorm is a widely used normalization technique in recent large language models [41]. Formally, RMSNorm takes two tensors, $X$ and $G$, as inputs and normalizes their element-wise products according to the root mean square:
$Y_{ij}=\frac{X_{ij}G_{j}}{\textrm{RMS}(X_{i})},\textrm{RMS}(X_{i})=\sqrt{\frac{1}{d}\sum_{j=1}^{d}X_{ij}^{2}},$ (1)